OpenStudy (anonymous):

An isosceles triangle has an area of 270 ft². If the base is 24 ft, what is the length of each leg?

OpenStudy (anonymous):



OpenStudy (anonymous):

the way i read the problem you know that \(270=\frac{1}{2}\times 24\times h\) making \(12h=270\) so \(h=\frac{45}{2}\)

OpenStudy (anonymous):

i am not sure what "height of each leg" means, i would say "length of each leg" and you can now find them via pythagoras